OpenStudy (anonymous):

Find the slope, if it exists, of the line containing the pair of points. (9,0) and ( 0,-5).

OpenStudy (anonymous):

-5/-9 = 5/9

OpenStudy (anonymous):

-5over9 Because you do this y2-y1 over x2-x1= your slope!

OpenStudy (anonymous):

So -5 over 9 would be the slope which is the answer.

OpenStudy (anonymous):

\[\frac{-5-0}{0 - 9}\]

OpenStudy (anonymous):

Slope is given by the formula: \[Slope = \frac{y_2 - y_1}{x_2 - x_1} \implies Slope = \frac{-5-0}{0-9} \implies \frac{-5}{-9} \implies \color{green}{Slope = \frac{5}{9}}\]

OpenStudy (anonymous):

5 over 9 :) yes is the slope!!!

OpenStudy (anonymous):

Thanks, guys.

OpenStudy (anonymous):

It will be more appreciable if you will remember the formula for slope and next time try at least once to find slope when you points are given to you.. That will increase your knowledge...

OpenStudy (anonymous):

*two instead of you..